OpenStudy (anonymous):

need help evaluating.. ∫sin(3x)cos(x)dx

OpenStudy (anonymous):

I know I have to split it up. But having problems doing so. I am not sure how to split up sin(3x), any ideas? I am sure this is an easy one I am just new

OpenStudy (anonymous):

no ideas?

OpenStudy (anonymous):

maybe you should start by making your odd number power even like this: \[\int\limits_{?}^{?}\sin^2(x)\cos^2(x)\sin(x) dx\]

OpenStudy (anonymous):

and then use trig substitution \[\sin^2x=1-\cos^2x\]

OpenStudy (dumbcow):

\[\sin(3x) = \sin(2x +x) = \sin(2x)\cos(x)+\sin(x)\cos(2x)\] \[\sin(2x) = 2\sin(x)\cos(x)\] \[\cos(2x) = 2\cos^{2}x -1\] \[\rightarrow \sin(3x) = 4\sin(x)\cos^{2}x - \sin(x)\] then use substitution u = cos x du = -sin x dx

OpenStudy (turingtest):

an easier way is to remember this all-too-forgotten trig identity:\[\sin\alpha\cos\beta=\frac12[(\sin(\alpha+\beta)+\sin(\alpha-\beta)]\]I don't know why they don't teach this very often...